### v2

Abstract
In the 1980s, Erdős and Sós initiated the study of Turán problems with a uniformity condition on the distribution of edges: the uniform Turán density of a hypergraph $H$ is the infimum over all $d$ for which any sufficiently large hypergraph with the property that all its linear-size subhypergraphs have density at least $d$ contains $H$. In particular, they asked to determine the uniform Turán densities of $K_{4}^{(3)-}$ and $K_{4}^{(3)}$. After more than 30 years, the former was solved in [Israel J. Math. 211 (2016), 349–366] and [J. Eur. Math. Soc. 20 (2018), 1139–1159], while the latter still remains open. Till today, there are known constructions of $3$-uniform hypergraphs with uniform Turán density equal to 0, $1/27$, $4/27$ and $1/4$ only. We extend this list by a fifth value: we prove an easy to verify sufficient condition for the uniform Turán density to be equal to $8/27$ and identify hypergraphs satisfying this condition.
1 Introduction
Turán problems, which ask for the minimum density threshold for the existence of a certain substructure, are one of the most fundamental kind of problems in extremal combinatorics. Formally, the Turán density of a graph $G$ is the infimum over all $d$ such that every sufficiently large $n$-vertex graph with $d\binom{n}{2}$ edges contains $G$ as a subgraph. The classical theorems of Mantel [16] and Turán [27] determine the Turán density of complete graphs. For general graphs, Erdős and Stone [9] proved that the Turán density of any $r$-chromatic graph is equal to $\frac{r-2}{r-1}$, see also [7]. While the situation is well-understood in the graph setting, the analogous questions in the hypergraph setting are among the most challenging problems in extremal combinatorics. Erdős [5] offered \$1 000 for determining the Turán density of all complete $k$-uniform hypergraphs for $k\geq 3$ and \$500 for determining the Turán density of any single complete $k$-uniform hypergraph (with at least $k+1$ vertices). Even determining the Turán density of $K_{4}^{(3)}$, the complete $3$-uniform hypergraph with $4$ vertices, has resisted all attempts for its resolution since its formulation 80 years ago [27]; we refer to [10, 3, 18] for partial and related results, and also to the surveys by Keevash [13] and Sidorenko [26].
Most of the known and conjectured extremal constructions for Turán problems have large independent sets, i.e., linear-size sets of vertices with no edges. This led Erdős and Sós [8, 6] to propose studying Turán problems with the additional requirement that the edges of the host hypergraph are distributed uniformly. Formally, the uniform Turán density of a $k$-uniform hypergraph $H$ is the infimum over all $d$ such that for every $\varepsilon>0$, there exists $n_{0}$ such that the following holds: every $k$-uniform hypergraph $H_{0}$ with $n\geq n_{0}$ vertices such that any subset of $n^{\prime}\geq\varepsilon n$ vertices of $H_{0}$ spans at least $d\binom{n^{\prime}}{k}$ edges contains $H$ as a subgraph. So, unlike in the case of Turán problems, uniform Turán problems require host hypergraphs to possess the required edge density on all linear-size vertex subsets. We remark that the notion is interesting for $k$-uniform hypergraphs with $k\geq 3$ only, since it can be shown that the uniform Turán density of every graph is equal to 0 (this follows e.g. from [25, Theorem 1]).
Erdős and Sós also asked to determine the uniform Turán density of the two smallest non-trivial $3$-uniform hypergraphs: the complete $3$-uniform hypergraph $K_{4}^{(3)}$ with four vertices and the $3$-uniform hypergraph $K_{4}^{(3)-}$, which is the hypergraph $K_{4}^{(3)}$ with an edge removed. Determining the uniform Turán density of $K_{4}^{(3)}$ remains a challenging open problem though it is believed that a 35-year-old construction of Rödl [25] showing that the uniform Turán density of $K_{4}^{(3)}$ is at least $1/2$ is optimal [19]. On a positive note, the uniform Turán density of $K_{4}^{(3)-}$ was shown to be equal to $1/4$ by Glebov, Volec and the third author [12] using arguments based on the flag algebra method of Razborov [17] and by Reiher, Rödl and Schacht [23] using direct combinatorial arguments.
In addition to $K_{4}^{(3)-}$, there are very restricted families of $3$-uniform hypergraphs whose uniform Turán density is known exactly. Reiher, Rödl and Schacht [21] classified $3$-uniform hypergraphs with uniform Turán density equal to 0. The first, third and fifth authors [11] constructed a family of $3$-uniform hypergraphs with uniform Turán density equal to $1/27$. Finally, Bucić, Cooper, Mohr, Munhá Correia and the third author [1] determined the uniform Turán density of $3$-uniform tight cycles of length at least five (the density is equal to $4/27$, if the length is not divisible by three, and it is 0 otherwise). Lastly, the approach from [23] was extended to a broader family of hypergraphs with uniform Turán density equal to $1/4$ [2, 15]. Hence there are only four values known to be the uniform Turán density of some hypergraph, and those are 0, $1/27$, $4/27$ and $1/4$. For further exposition, we refer the reader to the survey by Reiher [19] on uniform Turán densities of hypergraphs, which also includes results from e.g. [20, 24, 22] on stronger notions of uniform density.
Our main result is an easy to verify sufficient condition for a $3$-uniform hypergraph $H$ to have uniform Turán density equal to $8/27$. We then identify $3$-uniform hypergraphs $H$ that satisfy this condition, and so we add $8/27$ to the list of known uniform Turán densities. Since we mostly deal with $3$-uniform hypergraphs, for the rest of the paper we drop the adjective $3$-uniform when we talk about $3$-uniform hypergraphs except for those places where we wish to emphasize that a considered hypergraph is $3$-uniform in the interest of clarity (e.g., in the statements of lemmas and theorems).
To state our condition, we need the notion of a palette, which is inspired by the construction of Rödl from [25]; we also refer the reader to the survey [19] for further details on the notion. A palette is a set of ordered triples of colors. For example, the simplest palette is the palette $\Phi_{0}=\{(\alpha,\beta,\gamma)\}$, which consists of a single triple of mutually distinct colors denoted by $\alpha$, $\beta$ and $\gamma$. More complex examples of palettes are the palettes $\Phi_{3}$ and $\Phi_{8}$, which are defined below before the statement of Theorem 3. If $\Phi$ is a palette, we say that an $n$-vertex hypergraph $H$ is $\Phi$-colorable if there exists an ordering $v_{1},\ldots,v_{n}$ of the vertices of $H$ and a coloring of pairs of the vertices of $H$ with colors appearing in the palette $\Phi$ such that if the vertices $v_{i}$, $v_{j}$ and $v_{k}$, $1\leq i<j<k\leq n$, form an edge of $H$, then the triple $(c_{ij},c_{jk},c_{ik})$ is contained in $\Phi$ where $c_{ij}$ is the color of the pair $v_{i}$ and $v_{j}$, $c_{jk}$ is the color of the pair $v_{j}$ and $v_{k}$, and $c_{ik}$ is the color of the pair $v_{i}$ and $v_{k}$.
Having defined the notions of a palette and a $\Phi$-colorable hypergraph, we can state the above mentioned characterization of hypergraphs with the uniform Turán density equal to 0.
Theorem 1 (Reiher, Rödl and Schacht [21]).
The uniform Turán density of a $3$-uniform hypergraph $H$ is equal to 0 if and only if $H$ is $\Phi_{0}$-colorable.
Palettes also provide a method for lower bound constructions, which we present next; also see the survey [19]. We remark that a recent result of the last author [14], which is discussed in the concluding section, asserts that colorability by palettes fully determines uniform Turán densities. To present the lower bound construction method using palettes, we need an additional definition: the density of a palette $\Phi$, which consists of ordered triples of colors, is $|\Phi|/k^{3}$ where $k$ is the number of colors appearing in the triples of $\Phi$. For example, the density of the palette $\Phi_{0}$ is $1/27$.
Fix a palette $\Phi$ with density $d$ and an integer $n$. Let $k$ be the number of colors appearing in the triples of $\Phi$. We will construct an $n$-vertex hypergraph $H_{n}$ with vertices $v_{1},\ldots,v_{n}$ as follows. Color each pair of vertices uniformly at random with one of the $k$ colors appearing in the triples of $\Phi$ and include an edge formed by vertices $v_{i}$, $v_{j}$ and $v_{k}$, $1\leq i<j<k\leq n$, if the triple $(c_{ij},c_{jk},c_{ik})$ is contained in $\Phi$ where $c_{ij}$ is the color of the pair $v_{i}$ and $v_{j}$, $c_{jk}$ is the color of the pair $v_{j}$ and $v_{k}$, and $c_{ik}$ is the color of the pair $v_{i}$ and $v_{k}$. It is not hard to show that for every $\varepsilon>0$ and $\delta>0$, there exists an integer $n$ such that the following holds with positive probability: every subset of $n^{\prime}\geq\varepsilon n$ vertices of $H_{n}$ spans at least $(d-\delta)\binom{n^{\prime}}{3}$ edges. This construction yields that the uniform Turán density of any hypergraph that is not $\Phi$-colorable must be at least the density of $\Phi$. We state this conclusion as a proposition.
Proposition 2.
Let $\Phi$ be a palette. If a $3$-uniform hypergraph $H$ is not $\Phi$-colorable, then the uniform Turán density of $H$ is at least the density of $\Phi$.
Using Proposition 2 applied to the palette $\Phi_{0}$, we derive from Theorem 1 that the uniform Turán density of any hypergraph $H$ is either 0 or at least $1/27$ (recall that the construction of hypergraphs with uniform Turán density equal to $1/27$ was given in [11], so this jump value is the best possible).
To state our main result, we need the following two palettes, which we fix for the rest of the paper (the subscripts denote the number of triples contained in the palettes).
$\Phi_{3}$ $=\{(\alpha^{1},\beta^{1},\omega),(\alpha^{2},\omega,\gamma^{2}),(\omega,\beta^{3},\gamma^{3})\}$
$\Phi_{8}$ $=\{(x,y,z)\mbox{ such that }x\in\{\beta,\gamma\},y\in\{\alpha,\gamma\}\mbox{ and }z\in\{\alpha,\beta\}\;\}$
Note that the density of the palette $\Phi_{8}$ is $8/27$.
We are now ready to state our main result, which provides a condition that guarantees that the uniform Turán density of a hypergraph is equal to $8/27$.
Theorem 3.
The uniform Turán density of every $\Phi_{3}$-colorable $3$-uniform hypergraph that is not $\Phi_{8}$-colorable is $8/27$.
Since the density of the palette $\Phi_{8}$ is $8/27$, Theorem 3 follows from Proposition 2 and Theorem 10, which is proven in Section 3. In Section 4, we present a construction of hypergraphs that are $\Phi_{3}$-colorable but not $\Phi_{8}$-colorable, which establishes the existence of hypergraphs with uniform Turán density equal to $8/27$.
2 Preliminaries
In this section, we introduce the terminology related to partitioned hypergraphs. This framework, which is presented by Reiher in the survey [19], encapsulates hypergraph regularity arguments while avoiding some of the technicalities of a direct application of the Hypergraph Regularity Lemma.
An $N$-partitioned hypergraph $H$ is a $3$-uniform hypergraph such that its vertex set is partitioned to sets $V_{ij}$, $1\leq i<j\leq N$, and for every edge $e$ of $H$, there exist indices $1\leq i<j<k\leq N$ such that one vertex of $e$ is from $V_{ij}$, one from $V_{jk}$ and one from $V_{ik}$. An $(i,j,k)$-triad is the set of all edges of $H$ that have exactly one vertex in each of the sets $V_{ij}$, $V_{jk}$ and $V_{ik}$; note that each edge of $H$ belongs to exactly one triad. If $v$ is a vertex of $V_{ij}$, we write $d_{ij\to k}(v)$ for the number of edges of the $(i,j,k)$-triad that contain $v$ divided by $|V_{ik}|\cdot|V_{jk}|$; we use $d_{jk\to i}(v)$ and $d_{ik\to j}(v)$ analogously. Finally, the density of an $(i,j,k)$-triad is the number of edges forming the triad divided by $|V_{ij}|\cdot|V_{ik}|\cdot|V_{jk}|$, and the density of an $N$-partitioned hypergraph $H$ is the minimum density of a triad of $H$.
An $N$-partitioned hypergraph $H_{0}$ with parts $V_{ij}$, $1\leq i<j\leq N$, embeds an $n$-vertex hypergraph $H$ with vertices $v_{1},\ldots,v_{n}$ if it is possible to choose distinct $1\leq a_{1},\ldots,a_{n}\leq N$ corresponding to the vertices of $H$ and vertices $w_{ij}\in V_{a_{i}a_{j}}$ for $1\leq i<j\leq n$ such that if vertices $v_{i}$, $v_{j}$ and $v_{k}$ of $H$ form an edge, then $\{w_{ij},w_{jk},w_{ik}\}$ is an edge in the $(a_{i},a_{j},a_{k})$-triad of $H_{0}$. Reiher [19] provided a general statement that reduces proving an upper bound on the uniform Turán density of a hypergraph $H$ to embedding $H$ in partitioned hypergraphs.
Theorem 4 (Reiher [19, Theorem 3.3]).
Let $H$ be a $3$-uniform hypergraph and $d\in[0,1]$. Suppose that for every $\delta>0$ there exists $N$ such that every $N$-partitioned hypergraph with density at least $d+\delta$ embeds $H$. Then, the uniform Turán density of $H$ is at most $d$.
We remark that an extension of this general statement yields that the uniform Turán density of a hypergraph $H$ and any of its blowups are the same.
We next state four lemmas given in [1, Section 4], which we cast using our terminology; the lemmas were implicitly proven in [21] using a direct iterative approach, and alternative proofs based on Ramsey’s Theorem can be found in [11]. The lemmas are instances of the following metastatement: if a sufficiently large set $I_{0}\subseteq[N]$ is given and every triad of an $N$-partitioned hypergraph $H$ with indices from $I_{0}$ has a linear-size set of “good” vertices (for instance such a property might be having a high degree in the triad), then there is a large set of indices $I\subseteq I_{0}$ such that there is a universal choice of vertices that are good with respect to all triads with indices from $I$.
The first of the lemmas, which we prove for illustrative purposes, is the instance of the metastatement in the case when good vertices come from the set $V_{ik}$ of a $(i,j,k)$-triad.
Lemma 5 (Bucić, Cooper, Král’, Mohr and Munhá Correia [1, Lemma 4.3]).
For every $\varepsilon>0$ and $n$, there exists $n_{0}$ such that the following holds for every $N$-partitioned hypergraph $H$ with $N\geq n_{0}$, every $I_{0}\subseteq\{1,\ldots,N\}$ with $|I_{0}|\geq n_{0}$, and every choice of subsets $W_{ijk}\subseteq V_{ik}$, $i<j<k$, $i,j,k\in I_{0}$ such that $|W_{ijk}|\geq\varepsilon|V_{ik}|$: there exist $I\subseteq I_{0}$ with $|I|\geq n$ and $\gamma_{ik}$, $i<k$, $i,k\in I$, such that $\gamma_{ik}\in W_{ijk}$ for all $j\in I$ such that $i<j<k$.
Proof.
Consider $\varepsilon>0$ and $n$, and let $n_{0}$ be the Ramsey number such that every $2$-edge-colored $n$-uniform complete hypergraph with $n_{0}$ vertices contains a monochromatic complete hypergraph with $K=\max\{n^{2}-n+1,\lceil n/\varepsilon\rceil+2\}$ vertices. Fix an $N$-partitioned hypergraph $H$ with $N\geq n_{0}$, the set $I_{0}\subseteq\{1,\ldots,N\}$ such that $|I_{0}|\geq n_{0}$, and the subsets $W_{ijk}\subseteq V_{ik}$, $i<j<k$, $i,j,k\in I_{0}$ such that $|W_{ijk}|\geq\varepsilon|V_{ik}|$.
We now construct an auxiliary $2$-edge-colored $n$-uniform complete hypergraph with vertex set $I_{0}$ such that the edge formed by indices $i_{1}<\cdots<i_{n}$, $i_{1},\ldots,i_{n}\in I_{0}$, is colored blue if the sets $W_{i_{1},i_{k},i_{n}}$, $k=2,\ldots,n-1$, have a non-empty intersection, and it is colored red otherwise, i.e., the intersection of the sets $W_{i_{1},i_{k},i_{n}}$, $k=2,\ldots,n-1$, is the empty set. By the choice of $n_{0}$, there exists a $K$-element set $J\subseteq I_{0}$ such that all edges formed by the elements of $J$ have the same color. Let $i_{1}<\cdots<i_{K}$ be the elements of $J$.
First suppose that the common color of the edges formed by the elements of $J$ is red. By a simple averaging argument, the set $V_{i_{1},i_{K}}$ contains a vertex that is contained in at least $\frac{\varepsilon(K-2)|V_{i_{1},i_{K}}|}{|V_{i_{1},i_{K}}|}=\varepsilon(K-2)\geq n$ sets $W_{i_{1},i_{k},i_{K}}$, $k\in\{2,\ldots,K-1\}$; let $x$ be any such vertex of $V_{i_{1},i_{K}}$. However, any n2n-2 indices $i_{k}$, $k\in\{2,\ldots,K-1\}$, such that $x\in W_{i_{1},i_{k},i_{K}}$ together with $i_{1}$ and $i_{K}$ form a blue edge. Hence, the common color of the edges formed by the elements of $J$ cannot be red.
We next show that the set $I=\{i_{1},i_{n+1},i_{2n+1},\ldots,i_{(n-1)n+1}\}$ has the property from the statement of the lemma. To do so, we need to find elements $\gamma_{i_{a},i_{b}}$ for all $a,b\in\{1,n+1,\ldots,(n-1)n+1\}$ such that $a<b$. Fix such $i_{a}$ and $i_{b}$. Since the edge of the auxiliary complete hypergraph formed by the $(b-a)/n+1\leq n$ elements $i_{a},i_{a+n},\ldots,i_{b}$ and any n(ba)/n1n-(b-a)/n-1 elements among $a_{a+1},\ldots,a_{a+n-1}$ is blue, the sets $W_{i_{a},i_{k},i_{b}}$, $k=a+n,\ldots,b-n$, have a non-empty intersection, and we set $\gamma_{i_{a},i_{b}}$ to be any element contained in their intersection. ∎
The next two lemmas are the instance of the metastatement in the cases when good vertices come from the set $V_{ij}$ and the set $V_{jk}$ of a $(i,j,k)$-triad, respectively.
Lemma 6 (Bucić, Cooper, Král’, Mohr and Munhá Correia [1, Lemma 4.4]).
For every $\varepsilon>0$ and $n$, there exists $n_{0}$ such that the following holds for every $N$-partitioned hypergraph $H$ with $N\geq n_{0}$, every $I_{0}\subseteq\{1,\ldots,N\}$ with $|I_{0}|\geq n_{0}$, and every choice of subsets $W_{ijk}\subseteq V_{ij}$, $i<j<k$, $i,j,k\in I_{0}$ such that $|W_{ijk}|\geq\varepsilon|V_{ij}|$: there exist $I\subseteq I_{0}$ with $|I|\geq n$ and $\alpha_{ij}$, $i<j$, $i,j\in I$, such that $\alpha_{ij}\in W_{ijk}$ for all $k\in I$ such that $k>j$.
Lemma 7 (Bucić, Cooper, Král’, Mohr and Munhá Correia [1, Lemma 4.5]).
For every $\varepsilon>0$ and $n$, there exists $n_{0}$ such that the following holds for every $N$-partitioned hypergraph $H$ with $N\geq n_{0}$, every $I_{0}\subseteq\{1,\ldots,N\}$ with $|I_{0}|\geq n_{0}$, and every choice of subsets $W_{ijk}\subseteq V_{jk}$, $i<j<k$, $i,j,k\in I_{0}$ such that $|W_{ijk}|\geq\varepsilon|V_{jk}|$: there exist $I\subseteq I_{0}$ with $|I|\geq n$ and $\beta_{jk}$, $j<k$, $j,k\in I$, such that $\beta_{jk}\in W_{ijk}$ for all $i\in I$ such that $i<j$.
The last lemma is an extension to the setting when good vertices come from the set $V_{ij}$ of a $(a,i,j)$-triad, a $(i,b,j)$-triad and $(i,j,c)$-triad, i.e. which vertices are good depends on all three triads.
We conclude this section with another lemma proven in [1, Section 4].
Lemma 8 (Bucić, Cooper, Král’, Mohr and Munhá Correia [1, Lemma 4.2]).
For every $\varepsilon>0$ and $n$, there exists $n_{0}$ such that the following holds for every $N$-partitioned hypergraph $H$ with $N\geq n_{0}$, every $I_{0}\subseteq\{1,\ldots,N\}$ with $|I_{0}|\geq n_{0}$, and every choice of subsets $W_{ij}^{abc}\subseteq V_{ij}$ where $a,b,c,i,j\in I_{0}$ and $a<i<b<j<c$ such that $|W_{ij}^{abc}|\geq\varepsilon|V_{ij}|$: there exist $I\subseteq I_{0}$ with $|I|\geq n$ and $\omega_{ij}$, $i<j$, $i,j\in I$, such that $\omega_{ij}\in W_{ij}^{abc}$ for all $a,b,c\in I$ such that $a<i<b<j<c$.
3 Embedding
This section is devoted to showing that every sufficiently large hypergraph with uniform density larger than $8/27$ contains any fixed $\Phi_{3}$-colorable hypergraph $H$ as a subhypergraph. We start with the following simple lemma, which can be found as [11, Lemma 6]; we include a short proof for completeness. Recall that $d_{ij\to k}(v)$ is the number of edges of a $(i,j,k)$-triad that contain $v$ divided by $|V_{ik}|\cdot|V_{jk}|$, and $d_{jk\to i}(v)$ and $d_{ik\to j}(v)$ are used analogously.
Lemma 9.
For every $\varepsilon>0$ and $n$, there exists $n_{0}$ such that the following holds for every $N$-partitioned hypergraph $H$ with $N\geq n_{0}$: there exist $I\subseteq\{1,\ldots,N\}$ with $|I|\geq n$ and reals $a$, $b$ and $c$ such that the following holds for all $i,j,k\in I$ such that $i<j<k$:
$a$ $\leq\frac{|\{v\in V_{ij}\mbox{ with }d_{ij\to k}(v)\geq\varepsilon\}|}{|V_{ij}|}<a+\varepsilon,$
$b$ $\leq\frac{|\{v\in V_{jk}\mbox{ with }d_{jk\to i}(v)\geq\varepsilon\}|}{|V_{jk}|}<b+\varepsilon,\mbox{ and}$
$c$ $\leq\frac{|\{v\in V_{ik}\mbox{ with }d_{ik\to j}(v)\geq\varepsilon\}|}{|V_{ik}|}<c+\varepsilon.$
Proof.
Fix $\varepsilon\in(0,1)$ and $n$, and let $K=\lfloor\varepsilon^{-1}+1\rfloor^{3}$. Let $n_{0}$ be such that every $K$-edge-colored complete $3$-uniform hypergraph with $n_{0}$ vertices contains an $n$-vertex monochromatic complete hypergraph; such $n_{0}$ exists by Ramsey’s Theorem. Let $H$ be an $N$-partitioned hypergraph with $N\geq n_{0}$, and set
$s_{ij\to k}$ $=\frac{|\{v\in V_{ij}\mbox{ with }d_{ij\to k}(v)\geq\varepsilon\}|}{|V_{ij}|},$
$s_{jk\to i}$ $=\frac{|\{v\in V_{jk}\mbox{ with }d_{jk\to i}(v)\geq\varepsilon\}|}{|V_{jk}|},\mbox{ and}$
$s_{ik\to j}$ $=\frac{|\{v\in V_{ik}\mbox{ with }d_{ik\to j}(v)\geq\varepsilon\}|}{|V_{ik}|}.$
We next construct an auxiliary $K$-edge-colored complete $3$-uniform hypergraph $H^{\prime}$ with $N$ vertices: the color of the edge formed by the $i$-th, $j$-th and $k$-th vertex of $H^{\prime}$ is the triple $\left(\lfloor s_{ij\to k}/\varepsilon\rfloor,\lfloor s_{jk\to i}/\varepsilon\rfloor,\lfloor s_{ik\to j}/\varepsilon\rfloor\right)$. By the choice of $N$, there exist a subset $I\subseteq\{1,\ldots,N\}$ with $|I|=n$ and a triple $(a,b,c)$ such that all edges formed by the $i$-th, $j$-th and $k$-th vertex of $H^{\prime}$ have the color $(a,b,c)$ whenever $i,j,k\in I$. Since the set $I$ and the reals $a\cdot\varepsilon$, $b\cdot\varepsilon$ and $c\cdot\varepsilon$ have the properties given in the statement of the lemma, the proof of the lemma is finished. ∎
We are now ready to prove the main result of this section.
Theorem 10.
Let $H$ be a $3$-uniform hypergraph. If $H$ is $\Phi_{3}$-colorable, then the uniform Turán density of $H$ is at most $8/27$.
Proof.
Fix a $\Phi_{3}$-colorable hypergraph $H$ and let $n$ be the number of vertices of $H$. By Theorem 4, it is enough to show that for every $\delta>0$, there exists $N_{0}$ such that every $N_{0}$-partitioned hypergraph with density at least $8/27+\delta$ embeds $H$. Fix $\delta\in(0,1)$. We next choose values $N_{0},\ldots,N_{8}$ such that $N_{0}\gg N_{1}\gg\cdots\gg N_{7}\gg N_{8}$. To do so, set $\varepsilon=\delta/20$, $N_{8}=n$, and the values of $N_{0},\ldots,N_{7}$ as follows.
  • $N_{7}$ is the value of $n_{0}$ from Lemma 5 applied with $\varepsilon$ and $N_{8}$.
  • $N_{6}$ is the value of $n_{0}$ from Lemma 7 applied with $\varepsilon$ and $N_{7}$.
  • $N_{5}$ is the value of $n_{0}$ from Lemma 5 applied with $\varepsilon$ and $N_{6}$.
  • $N_{4}$ is the value of $n_{0}$ from Lemma 6 applied with $\varepsilon$ and $N_{5}$.
  • $N_{3}$ is the value of $n_{0}$ from Lemma 7 applied with $\varepsilon$ and $N_{4}$.
  • $N_{2}$ is the value of $n_{0}$ from Lemma 6 applied with $\varepsilon$ and $N_{3}$.
  • $N_{1}$ is the value of $n_{0}$ from Lemma 8 applied with $\varepsilon$ and $N_{2}$.
  • $N_{0}$ is the value of $n_{0}$ from Lemma 9 applied with $2\varepsilon$ and $N_{1}$.
For this value of $N_{0}$, we will show that every $N_{0}$-partitioned hypergraph with density at least $8/27+\delta$ embeds $H$.
Fix an $N_{0}$-partitioned hypergraph $H_{0}$ with density at least $8/27+\delta$. By Lemma 9, there exist $I_{1}\subseteq\{1,\ldots,N_{0}\}$ with $|I_{1}|\geq N_{1}$ and reals $a$, $b$ and $c$ such that
$a$ $\leq\frac{|\{v\in V_{ij}\mbox{ with }d_{ij\to k}(v)\geq 2\varepsilon\}|}{|V_{ij}|}<a+2\varepsilon,$
$b$ $\leq\frac{|\{v\in V_{jk}\mbox{ with }d_{jk\to i}(v)\geq 2\varepsilon\}|}{|V_{jk}|}<b+2\varepsilon,\mbox{ and}$
$c$ $\leq\frac{|\{v\in V_{ik}\mbox{ with }d_{ik\to j}(v)\geq 2\varepsilon\}|}{|V_{ik}|}<c+2\varepsilon,$
for all $i,j,k\in I_{1}$ and $i<j<k$.
We next show that $a+b+c$ is at least $2+3\varepsilon$, which we later show to imply that for every choice of $k<i<k^{\prime}<j<k^{\prime\prime}$ from $I_{1}$ the set $V_{ij}$ contains linearly many vertices $v$ such that $d_{ij\to k^{\prime\prime}}(v)\geq 2\varepsilon$, $d_{ij\to k}(v)\geq 2\varepsilon$, and $d_{ij\to k^{\prime}}(v)\geq 2\varepsilon$. Consider any of the triads, say the $(1,2,3)$-triad. Let $A$ be the set of all vertices $v\in V_{12}$ with $d_{12\to 3}(v)\geq 2\varepsilon$, $B$ the set of all vertices $v\in V_{23}$ with $d_{23\to 1}(v)\geq 2\varepsilon$, and $C$ the set of all vertices $v\in V_{13}$ with $d_{13\to 2}(v)\geq 2\varepsilon$. Since each vertex of $V_{12}\setminus A$ is in at most $2\varepsilon\cdot|V_{23}|\cdot|V_{13}|$ edges of the $(1,2,3)$-triad, each vertex of $V_{23}\setminus B$ is in at most $2\varepsilon\cdot|V_{12}|\cdot|V_{13}|$ edges of the $(1,2,3)$-triad, and each vertex of $V_{13}\setminus C$ is in at most $2\varepsilon\cdot|V_{12}|\cdot|V_{23}|$ edges of the $(1,2,3)$-triad, the $(1,2,3)$-triad has at most
$|A|\cdot|B|\cdot|C|+6\varepsilon\cdot|V_{12}|\cdot|V_{13}|\cdot|V_{23}|\leq\left(abc+12\varepsilon\right)\cdot|V_{12}|\cdot|V_{13}|\cdot|V_{23}|$
edges. Since the density of the $N_{0}$-partitioned hypergraph $H_{0}$ is at least $8/27+\delta$, if follows that $abc+12\varepsilon$ is at least $8/27+\delta=8/27+20\varepsilon$ and so $abc\geq 8/27+8\varepsilon$. Using the Inequality of Arithmetic and Geometric Means, we obtain that
$a+b+c\geq 3\sqrt[3]{abc}\geq 3\sqrt[3]{8/27+8\varepsilon}\geq 3(2/3+\varepsilon)=2+3\varepsilon.$
Our next step is to apply Lemma 8 to identify vertices $\omega_{ij}$ such that $d_{ij\to k^{\prime\prime}}(\omega_{ij})\geq 2\varepsilon$, $d_{ij\to k}(\omega_{ij})\geq 2\varepsilon$, and $d_{ij\to k^{\prime}}(\omega_{ij})\geq 2\varepsilon$ for every choice of $k<i<k^{\prime}<j<k^{\prime\prime}$ from a suitable index set $I_{2}$; the vertices $\omega_{ij}$ will correspond to the color $\omega$ from the palette $\Phi_{3}$. To apply the lemma, we define sets $W_{ij}^{kk^{\prime}k^{\prime\prime}}\subseteq V_{ij}$ for $i,j,k,k^{\prime},k^{\prime\prime}\in I_{1}$ such that $k<i<k^{\prime}<j<k^{\prime\prime}$. Consider such $k<i<k^{\prime}<j<k^{\prime\prime}$. Let $A$ be the set of all $v\in V_{ij}$ with $d_{ij\to k^{\prime\prime}}(v)\geq 2\varepsilon$, $B$ the set of all vertices $v\in V_{ij}$ with $d_{ij\to k}(v)\geq 2\varepsilon$, and $C$ the set of all vertices $v\in V_{ij}$ with $d_{ij\to k^{\prime}}(v)\geq 2\varepsilon$. Let $m_{d}$, $d\in\{0,1,2,3\}$, be the number of vertices of $V_{ij}$ contained in exactly $d$ sets among $A$, $B$ and $C$. Observe that
$(a+b+c)|V_{ij}|\leq|A|+|B|+|C|=m_{1}+2m_{2}+3m_{3}\leq 2(m_{1}+m_{2})+3m_{3}\leq 2|V_{ij}|+3m_{3}$
, which implies that there are at least $\frac{a+b+c-2}{3}|V_{ij}|\geq\varepsilon|V_{ij}|$ vertices of $V_{ij}$ contained in $A\cap B\cap C$. We set $W_{ij}^{kk^{\prime}k^{\prime\prime}}=A\cap B\cap C$. Lemma 8 implies that there exist $I_{2}\subseteq I_{1}$ with $|I_{2}|\geq N_{2}$ and $\omega_{ij}\in V_{ij}$, $i,j\in I_{2}$ and $i<j$, such that $\omega_{ij}\in W_{ij}^{kk^{\prime}k^{\prime\prime}}$ for all $i,j,k,k^{\prime},k^{\prime\prime}\in I_{2}$ such that $k<i<k^{\prime}<j<k^{\prime\prime}$. In particular, it holds that $d_{ij\to k}(\omega_{ij})\geq 2\varepsilon$ for all $\in I_{2}, i<ji<j$, and $k\in I_{2}\setminus\{i,j\}$ (regardless whether $k<i<j$, $i<k<j$ or $i<j<k$).
Our next goal is to identify vertices that will correspond to the colors $\alpha^{1}$ and $\beta^{1}$ of the palette $\Phi_{3}$. To identify those corresponding to the color $\alpha^{1}$, we will apply Lemma 6. Consider $i,j,k\in I_{2}$ such that $i<j<k$ and let $W_{ijk}$ be the set of all vertices $v\in V_{ij}$ contained together with $\omega_{ik}$ in at least $\varepsilon|V_{jk}|$ edges of the $(i,j,k)$-triad. Since the number of edges containing $\omega_{ik}$ in the triad $(i,j,k)$-triad is at most
$|W_{ijk}|\cdot|V_{jk}|+\varepsilon|V_{ij}\setminus W_{ijk}|\cdot|V_{jk}|\leq|W_{ijk}|\cdot|V_{jk}|+\varepsilon|V_{ij}|\cdot|V_{jk}|$
and $d_{ik\to j}(\omega_{ik})\geq 2\varepsilon$, we obtain that $W_{ijk}$ contains at least $\varepsilon|V_{ij}|$ vertices. Lemma 6 yields that there exist $I_{3}\subseteq I_{2}$ with $|I_{3}|\geq N_{3}$ and $\alpha^{1}_{ij}\in V_{ij}$, $i,j\in I_{3}$ and $i<j$, such that $\alpha^{1}_{ij}$ and $\omega_{ik}$ are contained together in at least $\varepsilon|V_{jk}|$ edges of the $(i,j,k)$-triad for all $i,j,k\in I_{3}$ such that $i<j<k$.
We next apply Lemma 7 to identify the vertices corresponding to the color $\beta^{1}$. For $i,j,k\in I_{3}$ such that $i<j<k$, we set $W_{ijk}$ to be the set of all vertices $v\in V_{jk}$ that form together with $\alpha^{1}_{ij}$ and $\omega_{ik}$ an edge of the $(i,j,k)$-triad; note that $W_{ijk}$ contains at least $\varepsilon|V_{jk}|$ vertices. Hence, Lemma 7 implies that there exist $I_{4}\subseteq I_{3}$ with $|I_{4}|\geq N_{4}$ and $\beta^{1}_{jk}\in V_{jk}$, $j,k\in I_{4}$ and $j<k$, such that $\alpha^{1}_{ij}$, $\beta^{1}_{jk}$ and $\omega_{ik}$ form an edge of the $(i,j,k)$-triad for all $i,j,k\in I_{4}$ such that $i<j<k$. In particular, the vertices $\alpha^{1}_{ij}$, $\beta^{1}_{jk}$ and $\omega_{ik}$ of the $(i,j,k)$-triad indeed correspond to the colors $\alpha^{1}$, $\beta^{1}$ and $\omega$ of the palette $\Phi_{3}$.
In the completely analogous way, we apply Lemmas 6 and 5 (in this order) to obtain $I_{6}\subseteq I_{4}$ with $|I_{6}|\geq N_{6}$, $\alpha^{2}_{ij}$, $i,j\in I_{6}$ and $i<j$, and $\gamma^{2}_{ik}$, $i,k\in I_{6}$ and $i<k$, such that $\alpha^{2}_{ij}$, $\omega_{jk}$ and $\gamma^{2}_{ik}$ form an edge of the $(i,j,k)$-triad for all $i,j,k\in I_{6}$ such that $i<j<k$, and so the vertices $\alpha^{2}_{ij}$ and $\beta^{2}_{ij}$ correspond to the colors $\alpha^{2}$ and $\beta^{2}$ of the palette $\Phi_{3}$. Then, we apply Lemmas 7 and 5 to obtain $I_{8}\subseteq I_{6}$ with $|I_{8}|\geq N_{8}$, $\beta^{3}_{jk}$, $j,k\in I_{6}$ and $j<k$, and $\gamma^{3}_{ik}$, $i,k\in I_{6}$ and $i<k$, such that $\omega_{ij}$, $\beta^{3}_{jk}$ and $\gamma^{3}_{ik}$ form an edge of the $(i,j,k)$-triad for all $i,j,k\in I_{8}$ such that $i<j<k$; again, the vertices $\alpha^{3}_{ij}$ and $\beta^{3}_{ij}$ correspond to the colors $\alpha^{3}$ and $\beta^{3}$ of the palette $\Phi_{3}$.
We now argue that the $N_{0}$-partitioned hypergraph $H_{0}$ embeds the hypergraph $H$. By the assumption of the theorem, the hypergraph $H$ is $\Phi_{3}$-colorable. Let $v_{1},\ldots,v_{n}$ be the vertices of $H$ listed in the order such that there exists a choice of $c_{ij}\in\{\omega,\alpha^{1},\beta^{1},\alpha^{2},\gamma^{2},\beta^{3},\gamma^{3}\}$, $i,j\in\{1,\ldots,n\}$ such that if vertices $v_{i}$, $v_{j}$ and $v_{k}$, $1\leq i<j<k\leq n$, form an edge of $H$, then $(c_{ij},c_{jk},c_{ik})\in\Phi_{3}$. Let $a_{i}$, $i\in\{1,\ldots,n\}$, be the $i$-th smallest index contained in $I_{8}$ (note that $I_{8}$ has at least $N_{8}=n$ elements). For $1\leq i<j\leq n$, set $w_{ij}$ among the vertices of $V_{ij}$ as follows:
[…]
The choice of the vertices $\omega_{ij}$, $\alpha^{1}_{ij}$, $\beta^{1}_{ij}$, $\alpha^{2}_{ij}$, $\gamma^{2}_{ij}$, $\beta^{3}_{ij}$ and $\gamma^{3}_{ij}$, $i,j\in I_{8}$ and $i<j$, yields that if vertices $v_{i}$, $v_{j}$ and $v_{k}$, $1\leq i<j<k\leq n$, form an edge of $H$, then the vertices $w_{ij}$, $w_{jk}$ and $w_{ik}$ form an edge of the $(a_{i},a_{j},a_{k})$-triad. We conclude that the $N_{0}$-partitioned hypergraph $H_{0}$ embeds $H$ and the proof of theorem is finished. ∎
4 Construction
In this section, we identify examples of $3$-uniform hypergraphs that are $\Phi_{3}$-colorable but not $\Phi_{8}$-colorable. The key step in the construction is the following proposition. Recall that a hypergraph is linear if any two edges have at most one vertex in common.
Proposition 11.
Suppose that there exists an $n$-vertex $m$-edge 55-uniform linear hypergraph such that $n!<(10/9)^{m}$. Then, there exists an $n$-vertex $(3m)$-edge 33-uniform hypergraph that is $\Phi_{3}$-colorable but not $\Phi_{8}$-colorable.
Let us first discuss which hypergraphs satisfy the assumption of Proposition 11. Observe that any edge-maximal $5$-uniform linear hypergraph with $n$ vertices has at least
$\frac{\binom{n}{5}}{10\binom{n-2}{3}}=\frac{n(n-1)}{200}=\Omega(n^{2})$
edges. It follows that if $n$ is sufficiently large, then any edge-maximal $5$-uniform linear hypergraph with $n$ vertices satisfies the assumption of Proposition 11. In particular, there exist constants $c<C$ such that for any $n$ and $m$ with $Cn\log n<m<cn^{2}$, there exists an $n$-vertex $m$-edge 55-uniform linear hypergraph satisfying the assumption of Proposition 11.
An explicit small example of a $5$-uniform hypergraph satisfying the assumption can be obtained from the $5$-dimensional affine space ${\mathbb{F}}_{5}^{5}$ of order $5$. Recall that the vertices of ${\mathbb{F}}_{5}^{5}$ are $5$-dimensional vectors over the $5$-element field ${\mathbb{F}}_{5}$ and any five-element set $\{a+x\cdot b,\;x\in{\mathbb{F}}_{5}\}$ where $a,b\in{\mathbb{F}}_{5}^{5}$ and $b\not=(0,0,0,0,0)$ is a line; note that each line corresponds to $5\cdot 4=20$ distinct choices of $a$ and $b$ and every pair of points is contained together in exactly one line. The vertex set of the sought $5$-uniform hypergraph is formed by the $n=5^{5}=3\,125$ points of the space and each of the $m=\frac{n(n-1)}{20}=\frac{1}{10}\binom{n}{2}=488\,125$ lines of the space forms an edge. It is easy to check that $n!<(10/9)^{m}$.
We now prove Proposition 11.
Proof of Proposition 11.
Fix an $n$-vertex $m$-edge 55-uniform linear hypergraph $H_{0}$ such that $n!<(10/9)^{m}$. We construct the $n$-vertex $(3m)$-edge 33-uniform hypergraph $H$ in a random way. The vertex set of $H$ is the same as that of $H_{0}$. In each edge $e$ of $H_{0}$ choose two vertices $v$ and $v^{\prime}$ randomly (independently of the other edges) and include to $H$ as edges the three triples containing $v$, $v^{\prime}$ and one of the three remaining vertices of $e$; see Figure 1 for an illustration of the construction. We now verify that the hypergraph $H$ is $\Phi_{3}$-colorable and that it is not $\Phi_{8}$-colorable with positive probability.
Figure 1: The illustration of the construction of the $3$-uniform hypergraph $H$ from the $5$-uniform hypergraph $H_{0}$.
We start by verifying that the hypergraph $H$ is $\Phi_{3}$-colorable. Fix any order $v_{1},\ldots,v_{n}$ of the vertices of $H$ (and so of $H_{0}$). If a pair of vertices $v_{i}$ and $v_{j}$ of $H_{0}$ is not contained in an edge of $H_{0}$, choose $c_{ij}$ arbitrarily. Otherwise, consider any edge $e$ of $H_{0}$, and let $v_{i}$ and $v_{j}$, $i<j$, be the two vertices chosen in the construction of $H$. Set $c_{ij}=\omega$. For each of the three remaining vertices of $e$, say $v_{k}$, set $c_{ik}=\alpha^{1}$ and $c_{kj}=\beta^{1}$ if $i<k<j$, set $c_{ki}=\alpha^{2}$ and $c_{kj}=\gamma^{2}$ if $k<i$, and set $c_{ik}=\beta^{3}$ and $c_{jk}=\gamma^{3}$ if $k>j$. Since the hypergraph $H$ is linear, the values of $c_{ij}$, $1\leq i<j\leq n$, are well-defined. It is straightforward to verify that if $\{v_{i},v_{j},v_{k}\}$, $1\leq i<j<k\leq n$, is an edge of $H$, then $(c_{ij},c_{jk},c_{ik})\in\Phi_{3}$. Hence, the hypergraph $H$ is $\Phi_{3}$-colorable.
We next show that the hypergraph $H$ is not $\Phi_{8}$-colorable with positive probability. Fix any order $v_{1},\ldots,v_{n}$ of the vertices of $H$. Since the $5$-uniform hypergraph $H_{0}$ is linear, the hypergraph $H$ is not $\Phi_{3}$-colorable with respect to the fixed order $v_{1},\ldots,v_{n}$ if and only if the $5$-uniform hypergraph $H_{0}$ has an edge $\{v_{a},v_{b},v_{c},v_{d},v_{e}\}$, $1\leq a<b<c<d<e\leq n$, such that the two chosen vertices in the construction of $H$ are the vertices $v_{b}$ and $v_{d}$. Indeed, if $H_{0}$ has such an edge, then there is no choice of a color $c_{bd}$ such that $(c_{ab},c_{bd},c_{ad})\in\Phi_{8}$, $(c_{bc},c_{cd},c_{bd})\in\Phi_{8}$, and $(c_{bd},c_{de},c_{be})\in\Phi_{8}$. And if $H_{0}$ has no such edge, we can choose the colors of the pairs of the vertices $v_{a},v_{b},v_{c},v_{d},v_{e}$ within each edge of $H_{0}$ and this choice does not affect the choices within other edges of $H_{0}$ as the hypergraph $H_{0}$ is linear. Hence, for any fixed order of the vertices $v_{1},\ldots,v_{n}$, the probability that there exists an assignment of colors to the pairs of vertices such that $(c_{ij},c_{jk},c_{ik})\in\Phi_{8}$ for every edge $\{v_{i},v_{j},v_{k}\}$, $1\leq i<j<k\leq n$, of $H_{0}$ is equal to $(9/10)^{m}$. The union bound implies that the probability that there exists a vertex ordering that admits such an assignment of colors to the pairs of vertices is at most $n!\cdot(9/10)^{m}<1$, i.e., the hypergraph $H$ is not $\Phi_{8}$-colorable with probability at least $1-n!\cdot(9/10)^{m}>0$. ∎
5 Conclusion
We are aware of several follow-up results concerning the uniform Turán densities of $3$-uniform hypergraphs, and we would like to briefly mention the one that we believe to be the most surprising. Proposition 2 asserts that if a hypergraph $H$ is not $\Phi$-colorable for a palette $\Phi$ with density $d$, then the uniform Turán density of $H$ is at least $d$. Since all known extremal constructions for uniform Turán densities are based on palette coloring constructions, see [19], it is natural to ask whether this is a general phenomenon. In the subsequent work [14], the fifth author shows that for every $\delta>0$ every sufficiently large partitioned hypergraph with density $d$ contains a partitioned subhypergraph with density $d-\delta$ that can be associated with a palette coloring construction. It follows that the uniform Turán density of any $3$-uniform hypergraph $H$ is equal to the supremum over all $d$ such that $H$ is not $\Phi$-colorable for some palette $\Phi$ with density $d$. This result brings a powerful tool for determining the uniform Turán densities of hypergraphs; for example, it was shown that every tight cycle of length at least five is $\Phi$-colorable for every palette $\Phi$ with density larger than $4/27$ by Cooper in 2018 [4], but the uniform Turán density of tight cycles was determined only 5 years later [1].